OpenStudy (anonymous):

hat is the value of the expression 3–4?

OpenStudy (anonymous):

what is the value of the expression 3–4?

OpenStudy (anonymous):

@cc_peach @XXKXX

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

1/81 is my answer for this...

OpenStudy (triciaal):

oh you mean 3^-4

OpenStudy (triciaal):

correct

OpenStudy (anonymous):

yaay!! swaggy <3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok how do u do this?

OpenStudy (triciaal):

you can shift the number 6 to get the symbol for exponent

OpenStudy (anonymous):

3^(-4)=1/3^4=1/81

OpenStudy (anonymous):

ok how does that work? like what does it look like

Nnesha (nnesha):

exponent rule \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] when you see negative exponent you should flip the fraction to make it positive exponent

OpenStudy (anonymous):

ok

Nnesha (nnesha):

that's how \[\huge\rm \frac{ x^{-m} }{ 1 }=\frac{ 1 }{ x^m }\]

OpenStudy (triciaal):